# GenAudit: Fixing Factual Errors in Language Model Outputs with Evidence

### v1

@@ A.2 Annotator Instructions
A.3 Inter-annotator agreement
@@
A.4 Derivation of baseline precision-recall trade-off
@@ A.5 Training details
Each training run was carried out on 2×2$\times$2 × Nvidia A6000 GPUs each of which has 49GB of memory. We used gradient accumulation and gradient checkpointing to reduce the peak memory requirements during training. The implementation was done using the Huggingface transformers (Wolf et al., 2019), Deepspeed (Rasley et al., 2020), Peft (Mangrulkar et al., 2022) and Bitsandbytes [^5] https://github.com/TimDettmers/bitsandbytes libraries.
A.6 Sample model outputs
We show some sample corrections suggested by GenAudit using the Flan-UL2 backend to summaries generated by LLMs in Figures 4567.
@@ A.6 Sample model outputs
GPT-4
You are provided a document and its summary. The summary may potentially contain facts which contradict with the document or are not supported by any evidence in the document. The last sentence of the summary is marked as a claim. Find and list sufficient sentences in the document to provide evidence for the claim. Make sure to provide evidence for all the supported facts in the claim. Then, revise the claim to remove or replace facts which are not supported by the document or are contradicted by it. Only make changes to the text of the claim when necessary. When you add new information to the claim, it must be only to fix a contradictory fact in the claim, and not for changing the style of the text.
GPT-3.5-turbo
You are provided a document and its summary. The summary may potentially contain facts which contradict with the document or are not supported by any evidence in the document. The last sentence of the summary is marked as a claim. Find and list sufficient sentences in the document to provide evidence for the claim, and then revise the claim to remove or replace facts which are not supported by the document or are contradicted by it. When you add new information to the claim, it must be only to fix a contradictory fact in the claim, and not for changing the style of the text.
Others
@@
Gemini-pro, GPT-3.5-turbo, GPT-4
Generate a summary for the following document in brief. When creating the summary, only use information that is present in the document. Generate the summary in free-form text without using bullet points.
Others
Generate a summary for the following document in brief. When creating the summary, only use information that is present in the document.

### v3

@@ B.2 Annotator Instructions (for Section 5)
B.3 Inter-annotator agreement (for Section 5)
@@
B.4 Creating data for human fact-checking performance evaluation (for Section 6)
We sampled news articles randomly from the XSum dataset to use as reference text, and used GPT-3.5-Turbo to generate their summaries. We further used GPT-3.5-Turbo to modify the summaries generated in this first round by adding factual errors. We used the prompt template used by Laban et al. (2023) to generate the modified summaries with added errors. [^7] https://github.com/salesforce/factualNLG/blob/master/prompts/summedits/step2_inconsistent.txt The prompt template is shown in Figure 4.
Document:
[DOCUMENT]
Consistent Summary:
[CONSISTENT_SUMMARY]
Given the document and consistent summary above, generate one slightly modified
version of the summary such that the modifications introduce a factual
inconsistency. For example, you can modify a number, date, or entity, and negate
or modify a statement. Here are some rules to follow:
- The modifications should change at most 5-6 words from the original summary,
 and keep the rest the same.
- The modifications should change  different parts of the original summary.
- Your modifications should be challenging to detect: modify minimally while
 still introducing a factual inconsistency.
- The factual inconsistency you introduce should be subtle. For example if you
 replace an entity, make sure you replace it with another entity from the document.
- The modified summary should start with "[FIRST_THREE_WORDS] [...]", and end
with "[LAST_THREE_WORDS]"
Figure 4: Prompt template used for introducing factual errors in summaries used for human fact-checking performance evaluation (Section 6). Taken from Laban et al. (2023)
B.5 Derivation of baseline precision-recall trade-off
@@ B.6 Training details
Each training run was carried out on 2×2$\times$2 × Nvidia A6000 GPUs each of which has 49GB of memory. We used gradient accumulation and gradient checkpointing to reduce the peak memory requirements during training. The duration of each training run varied depending on the model used but fell in the range of 30-70 hours. The implementation was done using the Huggingface transformers (Wolf et al., 2019), Deepspeed (Rasley et al., 2020), Peft (Mangrulkar et al., 2022) and Bitsandbytes [^8] https://github.com/TimDettmers/bitsandbytes libraries.
@@
Table 7: Links to datasets used in this work
@@
GPT-4 You are provided a document and its summary. The summary may potentially contain facts which contradict with the document or are not supported by any evidence in the document. The last sentence of the summary is marked as a claim. Find and list sufficient sentences in the document to provide evidence for the claim. Make sure to provide evidence for all the supported facts in the claim. Then, revise the claim to remove or replace facts which are not supported by the document or are contradicted by it. Only make changes to the text of the claim when necessary. When you add new information to the claim, it must be only to fix a contradictory fact in the claim, and not for changing the style of the text.
GPT-3.5-turbo You are provided a document and its summary. The summary may potentially contain facts which contradict with the document or are not supported by any evidence in the document. The last sentence of the summary is marked as a claim. Find and list sufficient sentences in the document to provide evidence for the claim, and then revise the claim to remove or replace facts which are not supported by the document or are contradicted by it. When you add new information to the claim, it must be only to fix a contradictory fact in the claim, and not for changing the style of the text.
Others You are provided a document and its summary. The summary may potentially contain factual errors. The last sentence of the summary is marked as a claim. Find all sentences in the document providing evidence for the claim, and then revise the claim to remove or replace unsupported facts.
@@
Gemini-pro, GPT-3.5-turbo, GPT-4 Generate a summary for the following document in brief. When creating the summary, only use information that is present in the document. Generate the summary in free-form text without using bullet points.
Others Generate a summary for the following document in brief. When creating the summary, only use information that is present in the document.